# Regularizing Learnable Feature Extraction for Automatic Speech Recognition

### v1

@@ 2 Related Work
Additionally, SpecAugment [8] has been shown to be particularly effective for ASR. It works by masking blocks of time frames or feature channels during training as well as performing time warping. Following its success, numerous variations have been proposed, e.g. dropping [17] or swapping [18] blocks instead of masking, masking based on phonemes or other semantic regions [19, 20] and more [21, 22, 23]. [24] proposes a variation of SpecAugment’s time warping for a wav2vec 2.0 model with learnable front-end. However, this modification does not appear to be specifically designed for learnable feature extraction and the experimental results do not isolate its individual contribution to the performance improvement.
@@ 3.2 Audio Perturbation
Nonlinear Amplitude Perturbation: The next perturbation involves a nonlinear distortion of the audio signal’s amplitudes. It is defined as:
@@ 4.1 Data
The Switchboard-1 Release 2 dataset [28] serves as training data for our experiments. The corpus contains of conversational English telephony speech sampled at . We use Hub5’00 as the development set and Hub5’01 for evaluation.
@@ 4.2 ASR Model
For our experiments, we use a connectionist temporal classification (CTC)-based ASR model [29]. We follow the setup from a previous work [30] and use a 4-gram language model (LM) during recognition. Feature extraction is done as explained in Section 3.1 including VGG-style subsampling, followed by dropout and a final linear layer producing representations with a shift of . Following results from preliminary experiments, standard preemphasis with is used for the supervised convolutional features, but not for log Mel. Also, for the log Mel baseline, it was beneficial to normalize the input waveform over time and the resulting features over the batch. Our encoder consists of 12 conformer [31] layers with a hidden dimension of 512. Compared to the original paper, we swap the position of convolution and multi-head self-attention modules. A final linear layer maps the encoder output to the target vocabulary, resulting in 74M parameters for the log Mel model. When using learnable features, there are around 21k parameters for the feature extractor and the number of parameters for the linear layer after VGG increases from 1.3M to 12.3M because of the almost 10 times larger feature dimension111The VGG output dimension is 32 times the feature dimension resulting in parameters for the linear layer. leading to 85M parameters in total. We train our model using the NAdam optimizer applying L2 weight decay [6] with a factor of 0.01 for the VGG frontend and 0.0001 for the encoder. Further, dropout [7] is applied with a probability of 0.1 for regularization. We split the data into 6 sub-epochs and train a total of 450 sub-epochs, evaluating our model on sub-epochs 400 to 450 in steps of 10 and select the best checkpoint based on word error rate (WER) on Hub5’00 for test evaluation. We apply SpecAugment from the first sub-epoch. To increase training stability, perturbations are only enabled after the first 4 full epochs by initializing the model with the corresponding checkpoint of the baseline. To keep the total number of training epochs constant, we stop training with perturbations 4 full epochs earlier. We use a one-cycle learning rate (OCLR) schedule starting from with a peak of after 180 sub-epochs going down to over another 180 sub-epochs and staying at that level for the final sub-epochs of training. The experiments can be run on a single consumer graphics processing unit (GPU) with VRAM (e.g. Nvidia RTX 3090) so that the barrier for reproduction is low. Also, the software is publicly accessible [32, 33] and the recipes are available online222https://github.com/rwth-i6/returnn-experiments/tree/master/2025-regularizing-learnable-features.
@@ 5.1 Audio Perturbation
For non-standard perturbations, selecting appropriate hyperparameters is more challenging due to the lack of prior references. To address this, we determined the minimum and maximum perturbation factors such that the audio still sounded natural based on the authors’ subjective perception. Based on initial experimental results, we further tuned the parameters.
@@ 5.3 Final Combination
To obtain the final best model, we combine the best SpecAugment variation with the best audio perturbations. We select tempo perturbation here as it performed best on both log Mel and learnable features. While we also tried combining different perturbations, this resulted in an overall worse performance, likely because the regularization effect is too strong in this case. Table 3 presents the results. As visible in the first three lines, tempo perturbation improves the log Mel baseline results on both Hub5’00 and Hub5’01. In combination with tempo perturbation, SpecAugment in the STFT-domain is not beneficial for log Mel features. In contrast, the combination yields the best results for the learnable front-end which are now within 0.1% absolute from the best log Mel results. We also observe that the gap between train and dev scores is significantly smaller than in Figure 1, showing that overfitting has been mitigated effectively. Therefore, we demonstrate that the presented regularization methods for training an ASR system with learnable features allow closing the performance gap to traditional features.

### v2

@@ 2 Related Work
Additionally, SpecAugment [8] has been shown to be particularly effective for ASR. It works by masking blocks of time frames or feature channels during training as well as performing time warping. Following its success, numerous variations have been proposed, e.g. dropping [17] or swapping [18] blocks instead of masking, masking based on phonemes or other semantic regions [19, 20] and more [21, 22, 23]. [24] proposes a variation of SpecAugment's time warping for a wav2vec 2.0 model with learnable front-end. However, this modification does not appear to be specifically designed for learnable feature extraction and the experimental results do not isolate its individual contribution to the performance improvement.
@@ 3.2 Audio Perturbation
Nonlinear Amplitude Perturbation: The next perturbation involves a nonlinear distortion of the audio signal's amplitudes. It is defined as:
@@ 4.1 Data
The Switchboard-1 Release 2 dataset [28] serves as training data for our experiments. The corpus contains of conversational English telephony speech sampled at . We use Hub5'00 as the development set and Hub5'01 for evaluation.
@@ 4.2 ASR Model
For our experiments, we use a connectionist temporal classification (CTC)-based ASR model [29]. We follow the setup from a previous work [30] and use a 4-gram language model (LM) during recognition. Feature extraction is done as explained in Section 3.1 including VGG-style subsampling, followed by dropout and a final linear layer producing representations with a shift of . Following results from preliminary experiments, standard preemphasis with is used for the supervised convolutional features, but not for log Mel. Also, for the log Mel baseline, it was beneficial to normalize the input waveform over time and the resulting features over the batch. Our encoder consists of 12 conformer [31] layers with a hidden dimension of 512. Compared to the original paper, we swap the position of convolution and multi-head self-attention modules. A final linear layer maps the encoder output to the target vocabulary, resulting in 74M parameters for the log Mel model. When using learnable features, there are around 21k parameters for the feature extractor and the number of parameters for the linear layer after VGG increases from 1.3M to 12.3M because of the almost 10 times larger feature dimension111The VGG output dimension is 32 times the feature dimension resulting in parameters for the linear layer. leading to 85M parameters in total. We train our model using the NAdam optimizer applying L2 weight decay [6] with a factor of 0.01 for the VGG frontend and 0.0001 for the encoder. Further, dropout [7] is applied with a probability of 0.1 for regularization. We split the data into 6 sub-epochs and train a total of 450 sub-epochs, evaluating our model on sub-epochs 400 to 450 in steps of 10 and select the best checkpoint based on word error rate (WER) on Hub5'00 for test evaluation. We apply SpecAugment from the first sub-epoch. To increase training stability, perturbations are only enabled after the first 4 full epochs by initializing the model with the corresponding checkpoint of the baseline. To keep the total number of training epochs constant, we stop training with perturbations 4 full epochs earlier. We use a one-cycle learning rate (OCLR) schedule starting from with a peak of after 180 sub-epochs going down to over another 180 sub-epochs and staying at that level for the final sub-epochs of training. The experiments can be run on a single consumer graphics processing unit (GPU) with VRAM (e.g. Nvidia RTX 3090) so that the barrier for reproduction is low. Also, the software is publicly accessible [32, 33] and the recipes are available online222https://github.com/rwth-i6/returnn-experiments/tree/master/2025-regularizing-learnable-features.
@@ 5.1 Audio Perturbation
For non-standard perturbations, selecting appropriate hyperparameters is more challenging due to the lack of prior references. To address this, we determined the minimum and maximum perturbation factors such that the audio still sounded natural based on the authors' subjective perception. Based on initial experimental results, we further tuned the parameters.
@@ 5.3 Final Combination
To obtain the final best model, we combine the best SpecAugment variation with the best audio perturbations. We select tempo perturbation here as it performed best on both log Mel and learnable features. While we also tried combining different perturbations, this resulted in an overall worse performance, likely because the regularization effect is too strong in this case. Table 3 presents the results. As visible in the first three lines, tempo perturbation improves the log Mel baseline results on both Hub5'00 and Hub5'01. In combination with tempo perturbation, SpecAugment in the STFT-domain is not beneficial for log Mel features. In contrast, the combination yields the best results for the learnable front-end which are now within 0.1% absolute from the best log Mel results. We also observe that the gap between train and dev scores is significantly smaller than in Figure 1, showing that overfitting has been mitigated effectively. Therefore, we demonstrate that the presented regularization methods for training an ASR system with learnable features allow closing the performance gap to traditional features.